# On the interaction of dilatancy and friction in the behavior of fluid–saturated sheared granular materials: a coupled Computational Fluid Dynamics–Discrete Element Method study

### v1

@@ Abstract
Frictional instabilities in fluid-saturated granular materials underlie critical natural hazards such as submarine landslides and earthquake initiation. Distinct failure behaviors emerge under subaerial and subaqueous conditions due to the coupled effects of mechanical deformation, interparticle friction, and fluid interactions. This study employs three-dimensional coupled Computational Fluid Dynamics–Discrete Element Method (CFD-DEM) simulations to investigate the collapse and runout dynamics of dense and loose granular assemblies across these settings. Parametric analyses reveal that pore pressure evolution plays a central role in governing failure mechanisms: dense assemblies stabilize through dilation, while loose assemblies undergo rapid compaction and fluidization, particularly under subaqueous conditions. Spatiotemporal analyses of coarse-grained fields further highlight strain-rate-dependent behavior driven by evolving porosity and effective stress. Both environments exhibit rate-strengthening behavior that scales with the inertial number ($I_{n}$) and viscous number ($I_{v}$), though driven by distinct mechanisms: subaerial systems are dominated by interparticle contact networks, whereas subaqueous systems are influenced by fluid drag, pore-pressure buildup, and lubrication. An analytical solution for excess pore pressure is compared with breaching-induced pressure distributions from CFD-DEM simulations, using input parameters derived from numerical triaxial DEM tests. The model captures fluid–particle coupling effectively, reproducing comparable excess pore pressures at steady state, while early-time discrepancies underscore the complexities of transient interactions. These findings advance the understanding of failure mechanics in saturated granular media and support the development of physics based models for mitigating hazards associated with subaqueous granular flows.
@@ I Introduction
Submarine slides and slope failures contribute significantly to the transport and redistribution of sediments from the continental shelf into the deep sea. This is a process that also often involves the formation of turbidity currents. Similar to subaerial landslides, submarine slides also have different modes of sliding that can create particular flow dynamics and different potentials for subsequent transport. These modes of failure include liquefaction, hydraulic cracking of sediment deposits, slip/sliding on a failure surface Pinet et al. (2008); Blum et al. (2010); Edwards, Lee, and Field (1995). An additional, more complex, and less studied mode of submarine flow is known as breaching flow slides or retrogressive slope failures. It is arguable that this mode of sliding may be a type of slow, gradual frictional slip and, in essence, a form of sliding on a failure surface. Field and experimental studies have demonstrated that submerged granular systems, as opposed to subaerial ones, experience distinct failure modes, including liquefaction slides, abrupt frictional slides, and slow breaching flow slides, driven by the interplay of mechanical, frictional, and fluid-mediated processes Rauter (2021); Montellà et al. (2023).
@@
A few recent experimental and continuum mechanics-based modeling studies by You et al. You, Flemings, and Mohrig (2012a, 2014a); You et al. (2014) illuminated the main processes that control the breaching phenomenon. They investigated the failure and sliding mechanics of steep subaqueous sediment deposits using laboratory flume experiments that were instrumented with pore pressure sensors and a sonar imaging system (to image the progress of breaching/sliding failure fronts). You et al.. You, Flemings, and Mohrig (2012a, 2014a); You et al. (2014) have shown that breaching takes place in dense fine-grained sediments (such as sandy silts) and that it is caused by a temporary build-up of negative pore pressure due to transient dilation of dense fine-grained sediments at the onset of sliding, the production of negative pore fluid pressure temporarily increases the effective stress and shear resistance of materials, which in turn arrests an otherwise developing failure surface. The subsequent dissipation of negative pore pressure results in a continuous process of slow sediment destabilization so that slow breaching continues. You, Flemings, and Mohrig (2014a) further showed there exists a spectrum of breaching type failure itself, which can be interpreted as a gradual transition from full breaching flow slides to dual-mode failure (alternate, cyclic changes between abrupt sliding and slow breaching), to abrupt frictional sliding, as the sediment deposit thickness increases for a given deposit length (in the primary direction of sliding). While some aspects of these processes have been demonstrated using continuum models of friction-fluid-coupled media, the underlying physical mechanisms, particularly the transient interplay between dilatancy and friction, remain poorly understood under earth’s near-surface stress conditions. Moreover, the mathematical and computational frameworks required to accurately model such coupled processes across scales are still underdeveloped and actively evolving Soulaine (2024).
@@
Dilatancy refers to the tendency of granular materials to change volume when subjected to shear deformation Reynolds (1885); Bolton (1986); Kabla and Senden (2009). This behavior plays a critical role in granular failure processes due to its direct coupling with pore-pressure evolution Goren et al. (2010). The resulting interaction governs transitions from stable creep to localized failure and strongly influences the mobility and run-out of landslides Finnegan and Saffer (2024). These transitions are often observed within shear zones, regions of concentrated deformation and elevated shear stress, which typically develop in response to externally applied loads or internal heterogeneities in the granular structure Dey et al. (2016). As deformation continues within these zones, changes in porosity and grain fabric promote frictional weakening Wan and Guo (1999), while pore-pressure feedbacks amplify instability by reducing the effective normal stress Deng et al. (2020). This coupling between mechanical and hydraulic responses creates conditions conducive to dynamic failure Iverson (2005). However, not all shear zones evolve toward instability; under certain conditions, dilatant hardening may increase the system’s resistance to slip, thereby stabilizing deformation. Such behavior has been linked to the occurrence of slow slip events, which are characterized by quasi-static rupture and recurrence intervals that match field observations Segall et al. (2010). The mechanical response of shear zones is further influenced by their initial density: dense zones typically undergo dilation during early deformation stages, with porosity eventually stabilizing after sufficient shear displacement Iverson (2005); Garagash and Rudnicki (2003). Environmental conditions, particularly saturation, also play a key role in modulating dilatant behavior. Under submerged conditions, the presence of pore fluid reduces interparticle friction Legros (2002), introduces buoyant uplift forces Cassar, Nicolas, and Pouliquen (2005), and lowers the effective weight of the material, thereby enhancing its potential for downslope mobility Vescovi, Marveggio, and Di Prisco (2020). These factors ultimately govern the flow dynamics of landslides, which may exhibit free-fall, inertial, or viscous regimes depending on the Stokes number ($S_{t}$) and the density ratio ($r$) Courrech du Pont et al. (2003).
@@
At the microscale, dense suspensions exhibit viscoplastic rheology governed by frictional and collisional interactions, evolving as a function of the viscous number, $I_{v}$ Boyer, Guazzelli, and Pouliquen (2011); Buettner, Guo, and Curtis (2020). Shear zone behavior is also modulated by shear rate Marone, Raleigh, and Scholz (1990), prevailing stress state Iverson et al. (2000); Aharonov and Sparks (2002), and the evolving balance between dilation and compaction. Collectively, these factors shape the stress–strain response and sliding stability of saturated granular systems Hsiao et al. (2017). In this context, the viscoinertial number offers a unified framework for characterizing rheological transitions in mono and bidisperse granular flows Cui et al. (2020).
@@
To understand the complex mechanisms driving granular failure, researchers integrate empirical observations with theoretical frameworks, laboratory experiments, and advanced numerical models. Laboratory studies confirm that porosity and dilatancy critically govern frictional strength, pore-pressure evolution, and landslide mobility. For instance, loose soils tend to liquefy abruptly under shear, whereas dense packings dilate, lowering pore pressure and resisting motion Iverson et al. (2000). Submerged granular collapse experiments reinforce these trends: loose beds produce rapid, shallow flows, while dense configurations yield slower, more coherent deposits due to enhanced fluid–grain coupling Rondon, Pouliquen, and Aussillous (2011). Classical theories such as Rowe’s stress-dilatancy framework offer insight into strength evolution under confinement but overlook key influences such as stress magnitude, void ratio, and initial packing conditions Wan and Guo (1999). To address these limitations, modern constitutive models incorporate state parameters and density-dependent dilatancy to better capture evolving material behavior.
@@
Advances in CFD-DEM modeling have further enabled high-resolution exploration of fluid–solid interactions under landslide-relevant conditions Pham et al. (2018). These simulations show that pore-pressure evolution governs shear localization: dilation induces suction and stabilizes dense packings, while compaction elevates pore pressure and promotes failure in looser configurations. Drainage and permeability conditions strongly modulate these responses, with dilation near drained boundaries increasing effective stress and compaction near impermeable zones reducing shear resistance Aharonov et al. (2013). CFD-DEM studies also reveal distinct flow regimes, from grain-inertial to fluid-inertial to viscous, each shaped by the Stokes number ($S_{t}$) and the density ratio ($r$), with important implications for landslide runout, energy dissipation, and deposition morphology Jing et al. (2019). Semi-resolved models further resolve vortex structures and fluid-mediated grain interactions, improving predictions of runout extent and final deposit structure Chen and Zhang (2023). Notably, bidisperse configurations, particularly those with coarse-over-fine layering, enhance segregation and kinetic energy transfer, promoting longer runouts, whereas reversed or randomly mixed beds tend to inhibit mobility Huang et al. (2025). Finally, the column aspect ratio influences vortex formation and segregation intensity; in submerged conditions, increased viscous drag reduces runout distances and delays recompaction Shademani, Blais, and Shakibaeinia (2021).
@@
Despite advances, the transient coupling of friction, pore pressure, and dilatancy in fluid-saturated granular systems, critical for flow initiation and localization, remains poorly understood. Hence, this study addresses these gaps using CFD-DEM simulations to investigate the roles of dilatancy, initial porosity, and frictional resistance in the shear deformation of dry and submerged granular systems. Systems with two initial packing densities are examined under subaerial and subaqueous conditions, incorporating Hertzian contact laws. We first detail the numerical model setup and boundary conditions, followed by a comparative analysis of failure kinematics and runout distances. We analyze spatiotemporal variations in coarse-grained stress, strain, and pore-pressure fields to identify emerging failure mechanisms and the evolution of shear zones. Through simulations, we explore how spatial dilation affects pore-pressure evolution and breaching in both dense and loose beds. We then characterize the evolution of friction, porosity, and pore pressure as functions of deviatoric strain rate and dimensionless numbers, including the inertial number and viscous number, offering a unified, strain-rate-dependent description of granular flow behavior across both dry and submerged conditions. A two-dimensional analytical pore-pressure model is employed later in the study to compare breaching erosion rates and pressure distributions. Triaxial DEM isotropic unloading tests are conducted to quantify isotropic compressibility and dilation strength. Dilation is modeled as a function of mean effective stress and the ratio of major to minor principal effective stresses.
@@ II Numerical approach : Coupled Computational Fluid Dynamics with the Discrete Element Method (CFD-DEM)
We use a numerical technique that combines Computational Fluid Dynamics (CFD) with the Discrete Element Method (DEM) to simulate fluid-particle systems. This approach enables detailed modeling of particle-fluid interactions by combining CFD’s ability to simulate fluid flows with DEM’s capability to track the motions and collisions of individual particles El-Emam et al. (2021). Below, we provide the details of the model setup and computations used in this study. We use the open-source code MFiX to simulate fluid–particle systems, which employs a hybrid Euler–Lagrange framework for multiphase flows, solving the fluid phase on an Eulerian grid and treating particles as discrete Lagrangian entities Garg et al. (2012); Musser et al. (2021). The CFD and DEM solvers are coupled through momentum exchange between the fluid and particles. Fluid motion is governed by the Navier–Stokes equations and is solved using a predictor–corrector method with an approximate projection scheme. Particle trajectories are determined by Newton’s laws, integrated using a forward Euler scheme with additional collision handling. Interpolation and deposition operators facilitate the coupling between the fluid and particle phases.
@@ II.1 Governing equations
The governing equations implemented in MFiX describe fluid-phase continuity and momentum conservation, excluding phase changes, chemical reactions, and phenomena such as growth, aggregation, or breakage Syamlal, Rogers, and OBrien (1993); Garg et al. (2012); Musser et al. (2021). Mass conservation, as expressed in Eq. 1, ensures continuity of the fluid phase by accounting for the volume fraction occupied by particles. Momentum conservation, given in Eq. 2, governs changes in fluid velocity resulting from pressure gradients, viscous stresses, body forces such as gravity, and interactions with particles. Fluid-phase properties are defined at cell centroids, while pressure is defined at cell nodes. The governing equations are solved using a predictor–corrector scheme with an approximate projection to enforce the divergence constraint. Key numerical methods include: second-order upwind discretization for convective terms, central differencing for viscous diffusion, and a nodal projection method to maintain incompressibility.
@@
The Lagrangian particle phase models individual particles and their interactions with both the surrounding fluid and neighboring particles. Each particle is characterized by its position, velocity, and angular velocity, as described by Eqs. 57. Particles are assumed to be spherical, with properties such as volume, mass, and moment of inertia computed from particle diameter and material density. Their motion is influenced by gravity, fluid–particle interaction forces (e.g., drag), and contact forces arising from collisions with other particles or walls. Particle collisions are modeled using a soft-sphere Hertzian contact model. Time integration is carried out using a forward Euler method, with subcycling applied during collision events. Particle volume is mapped to the Eulerian grid as a continuous volume fraction field using a transfer kernel to facilitate particle–fluid coupling. This Lagrangian–Eulerian approach enables detailed modeling of particle–particle and particle–fluid interactions, making it suitable for simulating complex multiphase flows. The total drag force acting on the $i^{\text{th}}$ particle within the $k^{\text{th}}$ cell and belonging to the $m^{\text{th}}$ solid phase is denoted by ${F}_{d}^{(i\in k,m)}$, and the net contact force from other particles is represented by ${F}_{c}^{(i)}$, both of which contribute to the linear momentum balance as shown in Eq. 6. The outward-pointing unit normal vector along the particle radius is denoted by $\eta$, and the total torque acting on the particle is ${T}^{(i)}$. The net force acting on a particle is ${F}_{T}^{(i)}$, which includes contributions from gravitational acceleration $g$, drag force $F_{d}$, and additional fluid-induced force $F_{f}$, as shown in Eq. 8. The fluid phase is governed by the Navier–Stokes equation as described by Eq. 9, which accounts for pressure gradients, viscous stresses, body forces such as gravity, and momentum exchange with particles. Within the Eulerian framework, fluid properties, including velocity and pressure, are defined on a fixed grid and evolve according to this equation. Fluid–particle coupling is implemented through source terms in the momentum equations, enabling two-way interactions necessary for capturing realistic multiphase dynamics. Drag forces are computed based on the relative velocity between particles and fluid.

### v2

@@ Abstract
Frictional instabilities in fluid-saturated granular materials underlie natural hazards, including submarine landslides and earthquake initiation. Experimental evidence shows distinct failure behaviors under subaerial and subaqueous conditions due to the coupled influences of mechanical deformation, interparticle friction, and particle–fluid interactions. We use three-dimensional coupled computational fluid dynamics–discrete element method (CFD–DEM) simulations to investigate the collapse and runout of dense and loose granular assemblies in both environments. Parametric analyses demonstrate that pore-pressure evolution controls the failure mode in saturated settings (fast vs slow sliding), consistent with prior laboratory experiments and lattice-Boltzmann–discrete element method simulations: dense assemblies stabilize via dilation, whereas loose assemblies compact rapidly and transiently fluidize. At the mesoscale, we coarse-grain particle-contact statistics and Eulerian fluid fields to define apparent friction and normalized pore pressure, and organize inertial and viscous responses using $\log_{10}(I_{n}/I_{v})$. Spatiotemporal analyses of these coarse-grained fields reveal strain-rate-dependent behavior governed by evolving porosity and effective stress. In both environments, friction in the failure shear zone is rate-strengthening with respect to the inertial number ($I_{n}$, for dry medium) and viscous number ($I_{v}$, for fluid-saturated medium). We further utilize the mesoscale stress framework to compare the evolution of pore pressure in the CFD–DEM simulations of subaqueous slope collapse with an analytical solution for the development of the failure front, using inputs derived from numerical triaxial DEM tests on the same assemblies. The analytical model reproduces steady-state excess pore pressures and captures fluid–particle coupling; however, a mismatch near failure onset suggests a role for transient frictional behavior in grain–fluid interactions. These results support the development of physics-based models of natural hazards and advance our mechanistic understanding of saturated granular failure.
@@ I Introduction
Submarine slides and slope failures play a major role in transporting and redistributing sediments from continental shelves to the deep sea, often accompanied by the formation of turbidity currents. As with subaerial landslides, submarine slides occur in multiple modes, each producing distinct flow dynamics and transport potentials. These modes include liquefaction, hydraulic cracking of sediment deposits, and slip along a failure surface pinet2008surface; blum2010assessing; edwards1995mudflow. A more complex and comparatively less studied mode is the breaching flow slide, also referred to as a retrogressive slope failure—a slow, gradual frictional slip along a failure surface you2014mechanics; You2012dynamics; eke2011field; van2002importance.
@@
Recent experimental and continuum mechanics–based modeling studies by You et al. you2014mechanics; You2012dynamics; you2014heterogeneity elucidate key processes governing breaching. Using laboratory flume experiments equipped with pore-pressure sensors and sonar imaging systems, they investigated the failure and sliding mechanics of steep subaqueous sediment deposits. Their results indicate that breaching occurs in dense, fine-grained sediments (e.g., sandy silts) due to a transient buildup of negative pore pressure during dilation at the onset of sliding. This negative pore pressure temporarily increases effective stress and shear resistance, arresting the formation of a continuous failure surface; as pore pressure dissipates, slow destabilization resumes, enabling progressive breaching. you2014mechanics further demonstrated that breaching spans a spectrum—from full breaching flow slides to dual-mode failures (cyclic alternation between abrupt sliding and slow breaching), and finally to abrupt frictional sliding—as deposit thickness increases for a given slope length. While aspects of these processes have been modeled using continuum approaches for friction–fluid coupled media, the underlying mechanisms—particularly the transient interplay between dilatancy and friction—remain poorly understood under Earth’s near-surface stress conditions. Mathematical and computational frameworks capable of capturing these coupled processes, ranging from discontinuous granular media to continuum scales, are an area of active research and development (Soulaine2024; montella2021two; rauter2021compressible; Guazzelli2018).
@@
Dilatancy—the tendency of granular materials to change volume under shear deformationReynolds1885; Bolton1986; Kabla2009 — plays a central role in granular failure through its direct coupling with pore-pressure evolution Goren2010. The interaction between dilatancy and pore pressure governs transitions from stable creep to localized failure and strongly influences landslide mobility and runout Finnegan2024. These transitions often occur within shear zones—regions of localized deformation and elevated shear stress that develop in response to external loads or internal structural heterogeneities Dey2016. As shear progresses, porosity and grain fabric evolve, promoting frictional weakening Wan1999, while pore-pressure feedbacks reduce effective normal stress and amplify instability Deng2020, creating conditions conducive to dynamic failure Iverson2005. Yet shear zones do not always evolve toward instability; under certain conditions, dilatant hardening can increase slip resistance and stabilize deformation, consistent with observations of slow slip events characterized by quasi-static rupture and field-scale recurrence intervals Segall2010. Shear-zone response also depends on initial density: dense zones typically dilate during initial shear deformation, with porosity stabilizing after sufficient shear displacement Iverson2005; Garagash2003. In submerged settings, pore fluid can reduce interparticle friction Legros2002, introduce buoyant uplift Cassar2005, and decrease the effective weight of the material, collectively enhancing downslope mobility Vescovi2020. These factors govern landslide flow dynamics, which may transition among free-fall, inertial, and viscous regimes depending on the Stokes number ($S_{t}$) and the density ratio ($r$) between the material and the ambient fluid CourrechduPont2003.
@@
At the microscale, dense suspensions exhibit a viscoplastic rheology governed by frictional and collisional interactions that evolve with the viscous number, $I_{v}$ Boyer2011; Buettner2020. Shear-zone behavior is further influenced by shear rate Marone1990, the prevailing stress state Iverson2000; Aharonov2002, and the evolving balance between dilation and compaction. Together, these factors define the stress–strain response and sliding stability of saturated granular systems Hsiao2017. Within this framework, the visco-inertial number offers a unified approach for characterizing rheological transitions in mono- and bidisperse granular flows Cui2020.
@@
Empirical observations, laboratory experiments, and numerical models collectively show that porosity and dilatancy critically influence frictional strength, pore-pressure evolution, and landslide mobility. Loose soils tend to liquefy abruptly under shear, whereas dense packings dilate, lowering pore pressure and resisting motion Iverson2000. Submerged granular-collapse experiments reinforce these trends: loose beds generate rapid, shallow flows, while dense configurations yield slower, more coherent deposits due to enhanced fluid–grain coupling Rondon2011. Classical theories, such as Rowe’s stress–dilatancy framework, provide valuable insights into strength evolution under confinement but overlook key parameters such as stress magnitude, void ratio, and initial packing conditions Wan1999. Modern constitutive models address these limitations by incorporating state parameters and density-dependent dilatancy to better represent evolving material behavior.
@@
Advances in CFD–DEM modeling have enabled high-resolution investigations of fluid–solid interactions under landslide-relevant conditions Pham2018. These simulations demonstrate that pore-pressure evolution controls shear localization: dilation induces suction and stabilizes dense packings, whereas compaction elevates pore pressure and promotes failure in looser configurations. Drainage and permeability strongly modulate these responses, with dilation near drained boundaries increasing effective stress and compaction near impermeable zones reducing shear resistance Aharonov2013. CFD–DEM studies also identify distinct flow regimes—grain-inertial, fluid-inertial, and viscous—governed by the Stokes number ($S_{t}$) and the density ratio ($r$), with important implications for runout, energy dissipation, and deposit morphology Jing2019. Semi-resolved models capture vortex structures and fluid-mediated grain interactions, improving predictions of runout extent and deposit architecture Chen2023. Bidisperse configurations, especially coarse-over-fine layering, enhance segregation and kinetic energy transfer, promoting longer runouts, whereas reversed or randomly mixed beds tend to inhibit mobility Huang2025. The Column aspect ratio also influences vortex formation and segregation intensity; under submerged conditions, increased viscous drag reduces runout distance and delays recompaction Shademani2021.
@@
Despite these advances, many of which correspond to the steady-state and final run-out distances of the sediment and granular deposits, the transient coupling among friction, pore pressure, and dilatancy in fluid-saturated granular systems—critical for flow initiation and localization—is not yet sufficiently explored and is not well-studied. This study addresses these gaps using CFD–DEM simulations to investigate the roles of dilatancy, initial porosity, and frictional resistance in the shear deformation of dry and submerged granular systems, with a focus on resolving the evolution of pore-fluid pressure and stress invariants within the failure zone. Two initial packing densities are examined under both subaerial and subaqueous conditions, incorporating Hertzian contact laws for grain–grain interactions. We first present the numerical model setup and boundary conditions, followed by a comparative analysis of failure kinematics and runout distances in the subaerial (dry) and subaqueous (fluid-saturated) cases. We then analyze spatiotemporal variations in coarse-grained stress, strain, and pore-pressure fields to identify emergent failure mechanisms and shear-zone evolution. The simulations are further used to examine how spatial dilation influences pore-pressure evolution and breaching-style failure in both dense and loose granular configurations. We characterize the evolution of friction, porosity, and pore pressure as functions of deviatoric strain rate and dimensionless numbers, including the inertial and viscous numbers. A two-dimensional analytical model for pore-pressure evolution is employed to compare theoretical predictions based on soil mechanics with the CFD–DEM results. To support this comparison, triaxial DEM isotropic unloading tests are conducted to quantify isotropic compressibility and dilation strength, with dilation modeled as a function of mean effective stress and the ratio of major to minor principal effective stresses.
This work advances studies of granular failure processes in fluid-saturated conditions by integrating simulation, analytical comparison, and mesoscale rheological characterization. Specifically, (i) we compare dilatancy–pressure coupling in subaqueous collapse through direct analysis of CFD–DEM–simulated pore-pressure fields and an analytical solution You2012dynamics, using DEM-derived elastic and poromechanical parameters ($E$ and $\beta$) You2012dynamics; you2014mechanics, thereby linking microstructural deformation to continuum-scale pore-pressure response in the context of breaching and slope-failure hazards Alhaddad2024-kj; (ii) we extract local rheological measures and unify inertial and viscous regimes within the combined $K$-framework Trulsson2012, $K=I_{v}+\alpha I_{n}^{2}$, enabling consistent comparison of subaerial and subaqueous failure behaviors; and (iii) we perform systematic micro-to-mesoscale upscaling from grain-resolved CFD–DEM data to Representative Volume Elements (RVEs) Rycroft2009, allowing coarse-grained definitions of apparent friction and normalized pore pressure that quantitatively relate evolving porosity, effective stress, and sliding dynamics. Collectively, these advances establish a mechanistic basis for understanding how transient fluid–grain interactions govern frictional instability and provide a framework for physics-based modeling of natural hazards involving fluid-saturated granular flows.
@@ II Numerical approach: Coupled Computational Fluid Dynamics with the Discrete Element Method (CFD-DEM)
We use a numerical technique that combines Computational Fluid Dynamics (CFD) with the Discrete Element Method (DEM) to simulate fluid-particle systems. This approach enables detailed modeling of particle-fluid interactions by combining CFD’s ability to simulate fluid flows with DEM’s capability to track the motion and collisions of individual particles ElEmam2021. Below, we provide details of the model setup and computations used in this study. We use the open-source code MFiX to simulate fluid–particle systems, which employs a hybrid Euler–Lagrange framework for multiphase flows, solving the fluid phase on an Eulerian grid and treating particles as discrete Lagrangian entities MFIXDEM2012; Musser2021. The CFD and DEM solvers are coupled through momentum exchange between the fluid and particles. Fluid motion is governed by the Navier–Stokes equations and is solved using a predictor–corrector method with an approximate projection scheme. Particle trajectories are determined by Newton’s laws, integrated using a forward Euler scheme with additional collision handling. Interpolation and deposition operators facilitate coupling between the fluid and particle phases.
@@ II.1 Governing equations
The governing equations implemented in MFiX describe fluid-phase continuity and momentum conservation, excluding phase changes, chemical reactions, and phenomena such as growth, aggregation, or breakage Syamlal1993; MFIXDEM2012; Musser2021. Mass conservation, as expressed in Eq. 1, ensures the continuity of the fluid phase by accounting for the volume fraction occupied by particles. Momentum conservation, given in Eq. 2, governs changes in fluid velocity resulting from pressure gradients, viscous stresses, body forces such as gravity, and interactions with particles. Fluid-phase properties are defined at cell centroids, while pressure is defined at cell nodes. The governing equations are solved using a predictor–corrector scheme with an approximate projection to enforce the divergence constraint. Key numerical methods include second-order upwind discretization for convective terms, central differencing for viscous diffusion, and a nodal projection method to maintain incompressibility.
@@
Set II Interphase coupling:
$\mathbf{I}_{fm}(\mathbf{x},t)=-\,\frac{1}{V_{c}}\sum_{p\in V_{c}}\mathbf{F}^{\,f\rightarrow p}_{\mathrm{hyd}}$ (3)
$\mathbf{F}^{\,f\rightarrow p}_{\mathrm{hyd}}=\mathbf{F}^{\mathrm{drag}}_{p}\;(+\,\mathbf{F}^{\mathrm{lift}}_{p}+\mathbf{F}^{\mathrm{vm}}_{p}\ \text{if enabled})$ (4)
Here, $V_{c}$ is the CFD cell volume and $\mathbf{I}_{fm}]={\rm N\,m^{-3}}$; if a projection kernel $W$ is used, replace $1/V_{c}$ by $W(\mathbf{x}-\mathbf{x}_{p})$ so that $\int_{V_{c}}\mathbf{I}_{fm}\,dV=-\sum_{p\in V_{c}}\mathbf{F}^{\,f\rightarrow p}_{\mathrm{hyd}}$. Action–reaction is enforced on the fluid control volume. In this work, we use hindered drag only (drag_type=’SYAM_OBRIEN’ in MFIX-DEM terminology MFIXDEM2012; MFIX_DEM_Ref); lift and virtual mass are disabled, while $\varepsilon_{f}\nabla p+\nabla\!\cdot(\varepsilon_{f}\bm{\tau}_{f})$ remain explicit in the fluid and $V_{p}\nabla p$ acts on particles Zhou2010JFM.
Hydrodynamic lubrication is the near-contact viscous resistance that grows sharply as interparticle gaps shrink (Davis1986). A scaling analysis (Carrara2019) suggests that explicitly accounting for lubrication forces is not required in our exact regime of saturated granular flows. Following the criteria discussed in Carrara et al. (2019)Carrara2019 and restated by Breard et al. (2022)Breard2022, the effects of lubrication forces become significant if both (i) $\log_{10}(X)>0$ with $X=Av_{p}/U_{T}$ and (ii) $\log_{10}(X)/\log_{10}(Y)>1$ with $Y=v_{f}/(2\alpha U_{T})$ and $\log_{10}(Y)>0$. Here, $A$ is a characteristic particle length, $v_{p}$ is the particle speed relative to the fluid, $U_{T}$ is the single-grain terminal settling velocity, $v_{f}$ is a characteristic fluid speed, and $\alpha$ is a dimensionless hindrance factor. We evaluated the parameters over $t\in[0,1.5]$ s. In the dense case ($\phi=0.579$) $\max\log_{10}(X)=-1.62$ (95th percentile 2.58-2.58). In the loose case ($\phi=0.444$) $\max\log_{10}(X)=-1.76$ (95th percentile 2.76-2.76). Thus, condition (i) never holds, so the joint criterion is never met, irrespective of the log–ratio values. This aligns with prior laboratory experiments Rondon2011 and two-phase flow modeling studies Pailha2009 that reproduced reduced subaqueous mobility and pore-pressure feedback without explicit lubrication.
The Saffman lift is a shear-induced transverse force on a small particle with finite slip in a viscous shear flow (Saffman1965). The classical Saffman lift law was derived for isolated particles (Saffman1965) and is incorporated into the general equation of motion for small rigid spheres in nonuniform flows (Maxey1983), but the lift coefficient is small, often orders of magnitude lower than drag for isolated spheres (Loth2008). Similarly, virtual mass effects become significant primarily under strong unsteady slip accelerations or in rapidly accelerating fluids (Auton1988), and are further illustrated in multiphase flow simulations (Ferrante2004). These conditions, however, are not representative of our column collapse simulation, where the governing assumptions and parameter ranges are explained in detail in Section II.2. In dense, viscous-dominated granular flows, two-phase simulations that include lift (with added mass and drag) report secondary influence relative to drag and pressure-gradient coupling (e.g., Chauchat2017SedFoam). Also, the reduced runout and pore-pressure feedback observed in laboratory experiments Rondon2011 and two-phase flow modeling studies Pailha2009 were captured using drag-only coupling, without using Saffman or virtual mass forces. Hence, we implemented hindered drag only, without Saffman lift or virtual mass, with our primary focus on comparing the rheology of saturated granular flows and the role of dilatancy in submerged granular collapses.
@@
The Lagrangian particle phase models individual particles and their interactions with both the surrounding fluid and neighboring particles. Each particle is characterized by its position, velocity, and angular velocity, as described by Eqs. 79. Particles are assumed to be spherical, with properties such as volume, mass, and moment of inertia computed from the particle diameter and material density. Their motion is influenced by gravity, fluid–particle interaction forces (e.g., drag), and contact forces arising from collisions with other particles or walls. Particle collisions are modeled using a soft-sphere Hertzian contact model. Time integration is carried out using a forward Euler method, with subcycling applied during collision events. Particle volume is mapped to the Eulerian grid as a continuous volume-fraction field using a transfer kernel to facilitate particle–fluid coupling. This Lagrangian–Eulerian approach enables detailed modeling of particle–particle and particle–fluid interactions, making it suitable for simulating complex multiphase flows. The total drag force acting on the $i^{\text{th}}$ particle within the $k^{\text{th}}$ cell and belonging to the $m^{\text{th}}$ solid phase is denoted by ${F}_{d}^{(i\in k,m)}$, and the net contact force from other particles is represented by ${F}_{c}^{(i)}$, both of which contribute to the linear momentum balance as shown in Eq. 8. The outward-pointing unit normal vector along the particle radius is denoted by $\eta$, and the total torque acting on the particle is ${T}^{(i)}$. The net force acting on a particle is ${F}_{T}^{(i)}$, which includes contributions from gravitational acceleration $g$, drag force $F_{d}$, and additional fluid-induced force $F_{f}$, as shown in Eq. 10. The fluid phase is governed by the Navier–Stokes equations as described by Eq. 11, which accounts for pressure gradients, viscous stresses, body forces such as gravity, and momentum exchange with particles. Within the Eulerian framework, fluid properties, including velocity and pressure, are defined on a fixed grid and evolve according to this equation. Fluid–particle coupling is implemented through source terms in the momentum equations, enabling two-way interactions necessary for capturing realistic multiphase dynamics. Drag forces are computed based on the relative velocity between particles and the fluid.